# Distilling the Implicit Multi-Branch Structure in LLMs’ Reasoning via Reinforcement Learning

### v3

@@ Abstract
Distilling reasoning paths from teacher to student models via supervised fine-tuning (SFT) provides a shortcut for improving the reasoning ability of the smaller Large Language Models (LLMs). However, the reasoning paths generated by teacher models often reflect only surface-level traces of their underlying authentic reasoning. Insights from cognitive neuroscience suggest that authentic reasoning involves a complex interweaving between meta-reasoning—which selects the appropriate sub-problem from multiple candidates—and solving, which addresses the sub-problem. It means that authentic reasoning has implicit multi-branch structure. Supervised fine-tuning collapses this rich structure into a flat sequence of token prediction in teacher’s reasoning path, which cannot distill this structure to student. To address this limitation, we propose RLKD, a reinforcement learning (RL)-based distillation framework guided by a novel Generative Structure Reward Model (GSRM). Our GSRM converts the reasoning path into multiple meta-reasoning-solving steps and gives the reward to measure the alignment between the reasoning structures of student and teacher. Our RLKD combines this reward with RL, enables the student LLM to internalize the teacher’s implicit multi-branch structure in authentic reasoning, rather than merely mimicking fixed teacher’s output paths. Experiments show that RLKD, even when trained on only 0.1% of the data under an RL-only regime, surpasses the performance of standard SFT-RL pipelines and further unleashes the potential reasoning ability of the student LLM than SFT-based distillation.
@@ 1 Introduction
Recently, Large Language Models (LLMs) have demonstrated impressive abilities on complex reasoning tasks Pan et al. (2025); Jiaqi et al. (2025); Xu et al. (2025a) via generating long reasoning paths Havrilla et al. (2024); Wei et al. (2022) such as Deepseek-R1 Guo et al. (2025). However, high training costs Guo et al. (2025); Xu et al. (2025b) and the strong base model Chu et al. (2025); Yue et al. (2025) are required in order for LLMs to emerge with this excellent capability, which prevents this reasoning capability from being explored by resource-constrained teams in developing their LLMs. To solve this challenge, supervised fine-tuning (SFT) on the reasoning paths generated from LLMs with powerful reasoning capabilities provides a shortcut and efficient method to make the smaller LLMs generate the long reasoning paths and achieve significant improvement Guo et al. (2025); Zhang et al. (2025); Face ; Wen et al. (2025). Despite this advance, some studies find SFT-distilled reasoning LLMs are trapped in rigid imitation rather than authentic reasoning. Purely mimicking the teacher’s reasoning paths can leave the student LLM “unthinking”: it replicates the surface form of the reasoning steps yet still makes errors on key underlying steps Chen et al. (2025); Dai et al. (2024).
To analyze and solve this phenomenon, we introduce the concepts from human cognitive neuroscience to rethink the definition of authentic reasoning of LLMs, which consists of two parts: one is meta-reasoning and the other is solving Stefik (1981); Cox and Raja (2007); Russell and Wefald (1991). Specifically, answering a complex problem involves multiple steps and each step consists of a meta-reasoning phase that determines the specific sub-problem from multiple potential sub-problems – followed by a solving phase that executes or answers the specific determined sub-problem. From this perspective, although the generated reasoning content is a definitive path, each step on this path is actually determined by meta-reasoning from multiple candidate states. Therefore, the complex interweaving between meta-reasoning and solving constitutes the authentic reasoning, in which both generated reasoning path and multiple other potential paths form the implicit multi-branch structure (Figure 1 (a)). A critical challenge in SFT-based distillation is that it does not teach the student LLM this structure. SFT trains a student LLM to imitate the teacher’s output sequence token-by-token with cross-entropy loss. Therefore, SFT collapses this rich implicit multi-branch structure in teacher into a flat sequence of token prediction to memorize only the teacher’s generated path (Figure 1 (b)) while fails to learn how to sample the path from other potential paths. Besides, by treating each token equally in training, the student LLM cannot learn how to meta-reason, how to solve and how to to connect them in multi-steps respectively.
@@
To solve this problem, it is important to provide the learning signal in distillation that clearly organizes multi-step meta-reasoning and solving in the reasoning paths, which is the core of implicit multi-branch structure. This is a highly semantic supervision that cannot be accomplished by token-level SFT and requires reinforcement learning (RL) Ouyang et al. (2022); Havrilla et al. (2024), so we propose RLKD, the first reinforcement learning-based knowledge distillation method for LLM’s reasoning. In RLKD, we design Generative Structure Reward Model (GSRM), a two-stage reward pattern that combines the semantic understanding of generative reward model Mahan et al. (2024) with the interpretability and controllability of the rule-based reward model Guo et al. (2025). GSRM can convert the reasoning path into the sequence consisting of multiple meta-reasoning-solving steps, and then score the matching degree between the sequences of the student and the teacher with structured reward mechanism. When combing GSRM with RL, our RLKD can guide the student LLM at a step-level on how to better perform sampling to select the most suitable sub-problem from multiple potential ones , and then solving it. In this way, RLKD can distill the implicit multi-branch structure in the reasoning of teacher LLMs to the student LLMs to make the student learn to perform authentic reasoning (Figure 1 (c)).
[…]
The contributions of this paper can be concluded as:
  • We point out that the reasoning paths generated by LLMs are merely superficial manifestations of authentic reasoning, whereas authentic reasoning possesses an implicit multi-branch structure, which can not be distilled to the student LLM by SFT.
  • We propose RLKD, which, to the best of our knowledge, is the first RL-based distillation method for LLM’s reasoning, and can effectively distill the implicit multi-branch structure.
  • The experimental results demonstrate the superiority of our proposed RL-based method over SFT-based distillation, highlighting the immense potential of RL-based distillation techniques in LLM’s reasoning and paving the way for new directions in subsequent research.
@@ 2.1 Structure in Reasoning of LLMs
Recent studies finds that the reasoning of LLMs has the structure. Chain-of-Thought Wei et al. (2022), Least-to-Most Zhou et al. (2022) and Self-Ask Press et al. (2022) initially formalize the reasoning path as a chain consisting of multiple nodes, which is a linear structure. Tree-of-Thought Yao et al. (2023), Graph-of-Thought Besta et al. (2024) and SearChain Xu et al. (2024a) explicitly builds a non-linear structure for reasoning. SuperCorrect Yang et al. (2025) uses high-level plans plus detailed steps as hierarchical thought templates to correct student models. The commonality of existing studies is that they explicitly let LLMs generated the specified structure in reasoning through prompt engineering or fine-tuning. Different from them, we introduce human cognitive neuroscience to rethink the reasoning, which consists of meta-reasoning and solving. The former is responsible for determining the sub-problem from multiple potential candidates, while the latter aims to solve the sub-problem. This means that a generated reasoning path contains multiple potential other paths, and these potential paths together with the generated path form the structure of reasoning. We focus on how to distill this implicit multi-branch structure into the student LLM.
2.2 Supervised Fine-Tuning for Reasoning Path Distillation
@@
Supervised fine-tuning (SFT) on chain-of-thought demonstrations has emerged as a straightforward way to distill reasoning capabilities from large models into smaller ones. For example, Deepseek releases a series distilled LLMs based on Deepseek-R1 and significantly improve the reasoning capabilities Guo et al. (2025). A similar study shows that with only  17k curated reasoning traces, a 32B student model can nearly match a closed-source o1-preview on math and coding benchmarks Li et al. (2025) and many open-source projects are released Wen et al. (2025); Face ; Bespoke-Labs . However, recent findings highlight that SFT often teaches format over substance: models learn to imitate the reasoning paths without authentic understanding of their content Chen et al. (2025); Dai et al. (2024). In fact, a student can produce correct answers by mimicking a long Chain-of-Thought (CoT) pattern even if many intermediate steps are incorrect Li et al. (2025). The key reason behind this limitation is that reasoning has implicit multi-branch structure but only SFT-based distillation collapses this rich structure into a flat sequence of token prediction. So we propose RLKD, a RL-based distillation method that can teach the student LLM to learn this structure.
2.3 Reinforcement Learning for Reasoning Optimization in LLMs
@@
Reinforcement learning (RL) has been explored as a means to optimize reasoning strategies in language models, building on foundations like Proximal Policy Optimization (PPO) Schulman et al. (2017). More recent advances include Group Relative Policy Optimization (GRPO) Shao et al. (2024), introduced by the DeepSeek team to push mathematical reasoning performance and Decoupled Clip and Dynamic Sampling Policy Optimization (DAPO) Yu et al. (2025) to train LLM with RL at Scale. These methods primarily focus on human feedback, outcome accuracy or heuristic rewards to optimize LLMs in specific downstream tasks. Different from them, our RLKD aims to use RL in knowledge distillation to transfer the implicit multi-branch structure from teacher to student to make the student LLM learn authentic reasoning.
@@ 3 Our Method
This section introduces details of our RLKD, a RL-based knowledge distillation method that can transfer the implicit multi-branch structure for complex reasoning from teacher to student. Firstly, we propose to train a Generative Structure Reward Model (GSRM) to score the alignment degree between the reasoning paths of student and teacher LLMs in terms of their implicit multi-branch structure. Then, we combine the GSRM with GRPO for RL-based knowledge distillation training.
@@ 3.1.1 Generation of Meta-Reasoning and Solving Sequence
We devise detailed instructions and examples to perform in-context learning (ICL) with the GPT-4o API, enabling it to automatically construct a large-scale supervised fine-tuning dataset. Each data sample of this dataset is a input-out pair: reasoning path $\mathbf{R}$ is the input and the sequence $\mathbf{S}$ consisting of multiple meta-reasoning and solving steps is the output. Specifically, we first structurally define each meta-reasoning and solving steps as ($\mathcal{M}$, $\mathcal{Q}$ $\mathcal{A}$), in which $\mathcal{M}$ is the content of meta-reasoning that focuses on determining the sub-problem that the current reasoning step should solve. $\mathcal{Q}$-$\mathcal{A}$ pair is the content of solving, in which $\mathcal{Q}$ is a clear description of the current sub-problem and $\mathcal{A}$ is the solving result to the sub-problem. So the sequence $\mathbf{S}$ containing $n$ meta-reasoning and solving steps can be described as:
$\mathbf{S}=[(\mathcal{M}_{1},\mathcal{Q}_{1}\&\mathcal{A}_{1}),(\mathcal{M}_{2},\mathcal{Q}_{2}\&\mathcal{A}_{2}),(\mathcal{M}_{3},\mathcal{Q}_{3}\&\mathcal{A}_{3}),...,(\mathcal{M}_{n},\mathcal{Q}_{n}\&\mathcal{A}_{n})].$ (1)
The essential for the generation of $\mathbf{S}$ is (1) containing each key reasoning step in the reasoning path and (2) fully decoupling the contents of $\mathcal{M}$, $\mathcal{Q}$, and $\mathcal{A}$ so that they contain and only contain the specified information. Based on these two points, we design specific instruction and examples covering four reasoning tasks (math, science, code and puzzles) to enable GPT-4o to perform effective ICL. The instruction is in Figure 2 (a) and one example is shown in Figure 2 (b), and the full prompt content for ICL can be found in Appendix.
@@
Besides, we introduce a verification-based feedback strategy to improve data quality. For each $\mathbf{S}$ generated by GPT-4o given input $\mathbf{R}$, we use Deepseek-V3 to determine whether $\mathbf{S}$ meets the requirements of the instruction. If it does not, we give the feedback from Deepseek-V3 to GPT-4o to re-generate $\mathbf{S}$. If a data sample fails to pass the verification after three re-generations, we discard this sample. We automatically execute this data production process on OpenThoughts-114k Team (2025), an open synthetic reasoning dataset with 114k high-quality examples covering math, science, code, and puzzles, and each example has Deepseek-R1 generated reasoning path, which is the $\mathbf{R}$. Finally, we get 93,62593,62593 , 625 $\mathbf{R}$-$\mathbf{S}$ pairs for the future supervised fine-tuning.
@@
This stage aims to train a generative reward model in our GSRM that can generate the sequence of meta-reasoning and solving pairs for the input reasoning path. Since the target output $\mathbf{S}$ (Equ. 1) is a highly structured text, we propose a training method called Structured Fine-grained Training to optimize each task (meta-reasoning generation and solving generation) in a fine-grained manner and dynamically adjust the optimization weight according to the difficulty of the task. Specifically, we split the tokens in $\mathbf{S}$ into two sets: one are tokens in meta-reasoning $\mathcal{M}$ and the other are tokens in solving $\mathcal{Q\&A}$. We perform 333 training epochs. Define $\mathcal{F}$ is the LLM used in this part (Qwen2.5-7B-Instruct), the first epoch is training for meta-reasoning generation, only loss from the tokens in $\mathcal{M}$ is computed as:
@@
The third stage is mixed training with dynamical weight. The total loss is the weighted average of the losses on the two token sets:
$\mathcal{L}_{3}=a\sum_{\mathbf{S}_{[i]}\in\mathcal{M}}-\log\mathcal{F}(\mathbf{S}_{[i]}|\mathbf{S}_{[1:i-1]};\theta)+b\sum_{\mathbf{S}_{[i]}\in\mathcal{Q\&A}}-\log\mathcal{F}(\mathbf{S}_{[i]}|\mathbf{S}_{[1:i-1]};\theta).$
$a$ and $b$ are initialized to the average of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ over training steps respectively. During training, we calculate the loss of the tokens in $\mathcal{M}$ and $\mathcal{Q\&A}$ respectively, and update $a$ and $b$ according to the ratio of the losses in two sets every $\alpha$ training steps. We set $\alpha$ as 100100100. Ablation study about this part can be found in Appendix.

### v4

# RLKD: Distilling LLMs’ Reasoning via Reinforcement Learning

@@ Abstract
Distilling reasoning paths from teacher to student models via supervised fine-tuning (SFT) provides a shortcut for improving the reasoning ability of the smaller Large Language Models (LLMs). However, the reasoning paths generated by teacher models often reflect only surface-level traces of their underlying authentic reasoning. Insights from cognitive neuroscience suggest that authentic reasoning involves a complex interweaving between meta-reasoning that selects the appropriate sub-problem from multiple candidates, and solving, which addresses the sub-problem. It means that authentic reasoning has implicit multi-branch structure. Supervised fine-tuning collapses this rich structure into a flat sequence of token prediction in teacher’s reasoning path, which cannot distill this structure to student. To address this limitation, we propose RLKD, a reinforcement learning (RL)-based distillation framework guided by a novel Generative Structure Reward Model (GSRM). Our GSRM converts the reasoning path into multiple meta-reasoning-solving steps and gives the reward to measure the alignment between the reasoning structures of student and teacher. Our RLKD combines this reward with RL, enables the student LLM to internalize the teacher’s implicit multi-branch structure in authentic reasoning, rather than merely mimicking fixed teacher’s output paths. Experiments show that RLKD, even when trained on only 0.1% of the data under an RL-only regime, surpasses the performance of standard SFT-RL pipelines and further unleashes the potential reasoning ability of the student LLM than SFT-based distillation. Code is available at {https://github.com/xsc1234/RLKD}.
@@ 1 Introduction
Recently, Large Language Models (LLMs) have demonstrated impressive abilities on complex reasoning tasks Pan et al. (2025); Jiaqi et al. (2025); Xu et al. (2025b) via generating long reasoning paths Havrilla et al. (2024); Wei et al. (2022) such as Deepseek-R1 Guo et al. (2025). However, high training costs Guo et al. (2025); Xu et al. (2025a) and the strong base model Chu et al. (2025); Yue et al. (2025) are required in order for LLMs to emerge with this excellent capability, which prevents this reasoning capability from being explored by resource-constrained teams in developing their LLMs. To solve this challenge, supervised fine-tuning (SFT) on the reasoning paths generated from LLMs with powerful reasoning capabilities provides a shortcut and efficient method to make the smaller LLMs generate the long reasoning paths and achieve significant improvement Guo et al. (2025); Zhang et al. (2025); Face ; Wen et al. (2025). Despite this advance, some studies find SFT-distilled reasoning LLMs are trapped in rigid imitation rather than authentic reasoning. Purely mimicking the teacher’s reasoning paths can leave the student LLM “unthinking”: it replicates the surface form of the reasoning steps yet still makes errors on key underlying steps Chen et al. (2025); Dai et al. (2024).
To analyze and solve this phenomenon, we introduce the concepts from human cognitive neuroscience to rethink the definition of authentic reasoning of LLMs, which consists of two parts: one is meta-reasoning and the other is solving Cox and Raja (2007); Russell and Wefald (1991). Specifically, answering a complex problem involves multiple steps and each step consists of a meta-reasoning phase that determines the specific sub-problem from multiple potential sub-problems – followed by a solving phase that executes or answers the specific determined sub-problem. From this perspective, although the generated reasoning content is a definitive path, each step on this path is actually determined by meta-reasoning from multiple candidate states. Therefore, the complex interweaving between meta-reasoning and solving constitutes the authentic reasoning, in which both generated reasoning path and multiple other potential paths form the implicit multi-branch structure (Figure 1 (a)). A critical challenge in SFT-based distillation is that it trains a student LLM to imitate the teacher’s output sequence token-by-token with cross-entropy loss. Therefore, SFT collapses this rich implicit multi-branch structure in teacher into a flat sequence of token prediction to memorize only the teacher’s generated path (Figure 1 (b)) while fails to learn how to sample the path from other potential paths.
@@
To solve this problem, it is important to provide the learning signal in distillation that clearly organizes multi-step meta-reasoning and solving in the reasoning paths, which is the core of implicit multi-branch structure. This is a highly semantic supervision that cannot be accomplished by token-level SFT and requires reinforcement learning (RL) Ouyang et al. (2022); Havrilla et al. (2024), so we propose RLKD, the first reinforcement learning-based knowledge distillation method for LLM’s reasoning. In RLKD, we design Generative Structure Reward Model (GSRM), a two-stage reward pattern that combines the semantic understanding of generative reward model Mahan et al. (2024) with the interpretability and controllability of the rule-based reward model Guo et al. (2025). GSRM can convert the reasoning path into the sequence consisting of multiple meta-reasoning-solving steps, and then score the matching degree between the sequences of the student and the teacher with structured reward mechanism. When combing GSRM with RL, our RLKD can guide the student LLM at a step-level on how to better perform sampling to select the most suitable sub-problem from multiple potential ones , and then solving it. In this way, RLKD can distill the implicit multi-branch structure in the reasoning of teacher LLMs to the student LLMs (Figure 1 (c)).
Experiments on math and graduate-Level Q&A show that our RLKD: (1) can use only $0.1\%$ training data with RL-only paradigm to outperform SFT-RL pipeline in Qwen2.5-Math, (2) can further unleash the potential reasoning ability of the student LLM than SFT-based distillation, and (3) can outperform existing RL baselines. In conclusion, this paper points out that the authentic reasoning possesses an implicit multi-branch structure, which can not be distilled to the student LLM by SFT and proposes RLKD, which is the first RL-based distillation method for LLM’s reasoning.
@@ 2.1 Structure in Reasoning of LLMs
Recent studies finds that the reasoning of LLMs has the structure. Chain-of-Thought Wei et al. (2022), Least-to-Most Zhou et al. (2022) and Self-Ask Press et al. (2022) initially formalize the reasoning path as a chain consisting of multiple nodes, which is a linear structure. Tree-of-Thought Yao et al. (2023), Graph-of-Thought Besta et al. (2024) and SearChain Xu et al. (2024a) explicitly builds a non-linear structure for reasoning. SuperCorrect Yang et al. (2025) uses high-level plans plus detailed steps as hierarchical thought templates to correct student models. The commonality of existing studies is that they explicitly let LLMs generated the specified structure in reasoning through prompt engineering or fine-tuning. We rethink the reasoning that a generated reasoning path contains multiple potential other paths, and these potential paths together with the generated path form the structure of reasoning. We focus on distilling this implicit multi-branch structure into the student LLM.
2.2 SFT for Reasoning Distillation
@@
Supervised fine-tuning (SFT) on chain-of-thought demonstrations has emerged as a straightforward way to distill reasoning capabilities from large models into smaller ones. For example, Deepseek releases a series distilled LLMs based on Deepseek-R1 and significantly improve the reasoning capabilities Guo et al. (2025). A similar study shows that with only  17k curated reasoning traces, a 32B student model can nearly match a closed-source o1-preview on math and coding benchmarks Li et al. (2025) and many open-source projects are released Wen et al. (2025); Face ; Bespoke-Labs . However, recent findings highlight that SFT often teaches format over substance: models learn to imitate the reasoning paths without authentic understanding of their content Chen et al. (2025); Dai et al. (2024). In fact, a student can produce correct answers by mimicking a long Chain-of-Thought (CoT) pattern even if many intermediate steps are incorrect Li et al. (2025). The key reason behind this is that reasoning has implicit multi-branch structure but SFT distillation collapses this structure into a flat sequence of token prediction. Our RL-based distillation can teach the student LLM to learn this.
2.3 Reinforcement Learning for LLMs’ Reasoning
@@
Reinforcement learning (RL) has been explored as a means to optimize reasoning strategies in language models, building on foundations like Proximal Policy Optimization (PPO) Schulman et al. (2017). More recent advances include Group Relative Policy Optimization (GRPO) Shao et al. (2024), introduced by the DeepSeek team to push mathematical reasoning performance and Decoupled Clip and Dynamic Sampling Policy Optimization (DAPO) Yu et al. (2025) to train LLM with RL at Scale. These methods primarily focus on human feedback, outcome accuracy or heuristic rewards to optimize LLMs in specific downstream tasks. Different from them, our RLKD aims to use RL in knowledge distillation.
@@ 3 Our Method
This section introduces details of our RLKD, a RL-based knowledge distillation method that can transfer the implicit multi-branch structure for complex reasoning from teacher to student. Firstly, we propose to train a Generative Structure Reward Model (GSRM) to score the alignment degree between the reasoning paths of student and teacher LLMs in terms of their implicit multi-branch structure. Then, we introduce the GSRM into RL-based knowledge distillation.
@@ 3.1 Generative Structure Reward Model
We devise detailed instructions and examples to perform in-context learning (ICL) with the GPT-4o API, enabling it to automatically construct a large-scale supervised fine-tuning dataset. Each data sample of this dataset is a input-out pair: reasoning path $\mathbf{R}$ is the input and the sequence $\mathbf{S}$ consisting of multiple meta-reasoning and solving steps is the output. Specifically, we first structurally define each meta-reasoning and solving steps as ($\mathcal{M}$, $\mathcal{Q}$ & $\mathcal{A}$), in which $\mathcal{M}$ is the content of meta-reasoning that focuses on determining the sub-problem that the current reasoning step should solve. $\mathcal{Q}$-$\mathcal{A}$ pair is the content of solving, in which $\mathcal{Q}$ is a clear description of the current sub-problem and $\mathcal{A}$ is the solving result to the sub-problem. So sequence $\mathbf{S}$ containing $n$ meta-reasoning and solving steps can be described as:
$\mathbf{S}=[(\mathcal{M}_{1},\mathcal{Q}_{1}\&\mathcal{A}_{1}),(\mathcal{M}_{2},\mathcal{Q}_{2}\&\mathcal{A}_{2}),...,(\mathcal{M}_{n},\mathcal{Q}_{n}\&\mathcal{A}_{n})].$
The essential for the generation of $\mathbf{S}$ is (1) containing each key reasoning step in the reasoning path and (2) fully decoupling the contents of $\mathcal{M}$, $\mathcal{Q}$, and $\mathcal{A}$ so that they contain and only contain the specified information. Based on these two points, we design specific instruction and examples covering four reasoning tasks (math, science, code and puzzles) to enable GPT-4o to perform effective ICL. One example is shown in Figure 2 (b).
@@
Besides, we introduce a verification-based feedback strategy to improve data quality. For each $\mathbf{S}$ generated by GPT-4o given input $\mathbf{R}$, we use Deepseek-V3 to determine whether $\mathbf{S}$ meets the requirements of the instruction. If it does not, we give the feedback from Deepseek-V3 to GPT-4o to re-generate $\mathbf{S}$. If a data sample fails to pass the verification after three re-generations, we discard this sample. We automatically execute this data production process on OpenThoughts-114k Team (2025), an open synthetic reasoning dataset with 114k high-quality examples covering math, science, code, and puzzles, and each example has Deepseek-R1 generated reasoning path, which is the $\mathbf{R}$. Finally, we get $93,625$ $\mathbf{R}$-$\mathbf{S}$ pairs for the future supervised fine-tuning.
@@
This stage aims to train a generative reward model in our GSRM that can generate the sequence of meta-reasoning and solving pairs for the input reasoning path. Since the target output $\mathbf{S}$ is a highly structured text, we propose a training method called Structured Fine-grained Training to optimize each task (meta-reasoning generation and solving generation) in a fine-grained manner and dynamically adjust the optimization weight according to the difficulty of the task. Specifically, we split the tokens in $\mathbf{S}$ into two sets: one are tokens in meta-reasoning $\mathcal{M}$ and the other are tokens in solving $\mathcal{Q\&A}$. We perform $3$ training epochs. $\mathcal{F}$ is the LLM used in this part (Qwen2.5-7B-Instruct), the first epoch is training for meta-reasoning, only loss from the tokens in $\mathcal{M}$ is considered:
@@
The third epoch is mixed training. Total loss is the dynamically weighted average of the losses on the two token sets:
$\mathcal{L}_{3}$ $=a\sum_{\mathbf{S}_{[i]}\in\mathcal{M}}-\log\mathcal{F}(\mathbf{S}_{[i]}|\mathbf{S}_{[1:i-1]};\theta)$ (1)
$+b\sum_{\mathbf{S}_{[i]}\in\mathcal{Q\&A}}-\log\mathcal{F}(\mathbf{S}_{[i]}|\mathbf{S}_{[1:i-1]};\theta).$
$a$ and $b$ are initialized to the average of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ over training steps respectively. During training, we calculate the loss of the tokens in $\mathcal{M}$ and $\mathcal{Q\&A}$ respectively, and update $a$ and $b$ according to the ratio of the losses in two sets every $\alpha$ training steps. We set $\alpha$ as $100$.